OpenStudy (anonymous):

Sin(theta) = 0.25 {(Theta): 0 < (theta) < pi} How do I solve it over that domain?

OpenStudy (campbell_st):

if \[\sin(\theta) = 0.25... then .... \theta = \sin^{-1}(0.25)\] your answer will need to be in radians and there are 2 angles 1st quadrant and 2nd quadrant where you need to use \[\pi - \theta\]

OpenStudy (anonymous):

Thank you :)